Sally and Bill work at different jobs. Sally earns $12 per hour and Bill earns $10 per hour. They each earn the same amount per week but Bill works 3 more hours. How many hours a week does Bill work?
Answer by richwmiller(17219)

12x=10(x+3)
12x=10x+30
2x=30
x=15 for Sally
x+3=18 hours for Bill
check
12*15=10*18
2*6*3*5=2*5*3*6
You can see the same factors.
ok